OpenStudy (republican31):

a. You have a 10 volt parallel circuit, with 2 resistors on it. What is the voltage across the first resistor? Across the second? b. You have the same two resistors on a 10 volt series circuit. Will the voltage going into the second resistor be more, less, or the same as that going into the first resistor? Exact numbers aren’t needed! c. You have a series circuit with a current of 6 amps and three resistors on it, with resistances of 10 ohms, 5 ohms, and 6 ohms, respectively. What is the voltage of this circuit? Show your calculation.

OpenStudy (anonymous):

Resistors in parallel \[\frac{ 1 }{ R _{total} }=\frac{ 1 }{ R _{1} }+\frac{ 1 }{ R _{2} }+...\] Resistors in series \[R _{total} = R _{1}+R _{2}+...\] V = I R

OpenStudy (radar):